# Parameter-Efficient Fine-Tuning in Spectral Domain for Point Cloud Learning

### v1

@@ 2.3 Spectral Methods for Point Clouds
Methods that utilize spectral information to understand point clouds achieve notable advancements [66, 5, 30, 43]. For example, Wang et al. [71] combine spectral feature learning with recursive clustering to overcome the isolated local feature aggregation. Zhang et al. [95] introduce tensor-based methods to estimate hypergraph spectrum components and frequency coefficients of point clouds in both ideal and noisy settings. GSDA [39] proposes the graph spectral domain attack, perturbing graph transform coefficients in the spectral domain. PointWavelet [76] designs graph wavelet transform to avoid the very time-consuming spectral decomposition. Ramasinghe et al. [54] propose a spectral-domain GAN that generates novel 3D shapes by using a spherical harmonics-based representation.
In this paper, we focus on the practical Parameter-efficient Fine-Tuning (PEFT) task, and propose a fresh PEFT method named PointGST for point cloud tasks, which efficiently fine-tunes the given pre-trained point cloud models in the spectral domain. The fundamental goals and techniques between our method and previous spectral-based point cloud methods are different.
@@ 3 Preliminary
This section revisits the mainstream fine-tuning paradigms and introduces the concept of Graph Fourier Transform.
3.1 Fine-Tuning Paradigm
Fully Fine-Tuning (FFT) is the most commonly used fine-tuning paradigm. Given a downstream training set $\boldsymbol{\Gamma}\left(\boldsymbol{x};\boldsymbol{y}\right)$ ($\boldsymbol{x}$ is the input and $\boldsymbol{y}$ is the label), the FFT trains the network $\mathcal{F}\left(\boldsymbol{x};\boldsymbol{\theta}\right)$ and obtains the updated weight $\boldsymbol{\theta}^{*}$, where $\boldsymbol{\theta}$ is the trainable parameters of $\mathcal{F}$. Specifically, the FFT paradigm can be formulated as:
$\boldsymbol{\theta}^{*}=\mathop{\arg\min}\limits_{\boldsymbol{\theta}}\ell\left(\mathcal{F}\left(\boldsymbol{x};\boldsymbol{\theta}\right),\boldsymbol{y}\right),$ (1)
where $\ell$ represents the loss function for downstream tasks. Notably, the FFT updates $\boldsymbol{\theta}$ to $\boldsymbol{\theta}^{*}$, and $\boldsymbol{\theta}^{*}$ retains the same shape as $\boldsymbol{\theta}$ after fine-tuning. In general, FFT will bring notable training costs reflected in GPU memory and storage.
Parameter-Efficient Fine-Tuning (PEFT) provides a practical solution by adapting pre-trained models to downstream tasks in a more efficient manner. Existing PEFT methods focus on tuning only a small subset of parameters $\boldsymbol{\theta}^{\prime}$. The optimized parameter $\boldsymbol{\theta}^{\prime*}$ can be represented as:
$\boldsymbol{\theta}^{\prime*}=\mathop{\arg\min}\limits_{\boldsymbol{\theta}^{\prime}}\ell\left(\mathcal{F}\left(\boldsymbol{x};\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right),\boldsymbol{y}\right),\left|\boldsymbol{\theta}^{\prime}\right|\ll\left|\boldsymbol{\theta}\right|.$ (2)
The magnitude of $\left|\boldsymbol{\theta}^{\prime*}\right|$ equals $\left|\boldsymbol{\theta}^{\prime}\right|$ and is significantly smaller than the pre-trained $\boldsymbol{\theta}$. During fine-tuning, the pre-trained $\boldsymbol{\theta}$ remains fixed while only $\boldsymbol{\theta}^{\prime}$ is update to $\boldsymbol{\theta}^{\prime*}$. Techniques for incorporating $\boldsymbol{\theta}^{\prime}$ include adding lightweight additional parameters and selecting a small subset of the pre-trained $\boldsymbol{\theta}$.
@@ 3.2 Graph Fourier Transform
Let us define a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\boldsymbol{\mathcal{W}}\}$, where $\mathcal{G}$ is the set of vertices ($|\mathcal{G}|=n$), $\mathcal{E}$ represents the edges, and $\boldsymbol{\mathcal{W}}$ is the adjacency matrix. In our context, a graph signal $\boldsymbol{z}\in\mathbb{R}^{n\times c}$ assigns a vector ($\boldsymbol{z}_{i}\in\mathbb{R}^{1\times c}$) to each vertex, while the entry $w_{i,j}$ in $\boldsymbol{\mathcal{W}}$ indicates the weight of the edge between vertices $i$ and $j$, reflecting their similarity. Then the Laplacian matrix is defined as $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{\mathcal{W}}$ [59], where $\boldsymbol{D}$ is a diagonal matrix with each element $d_{i,i}=\sum_{j=0}^{n-1}w_{i,j}$ representing the degree of vertex $i$. For an undirected graph with non-negative weights, $\boldsymbol{L}\in\mathbb{R}^{n\times n}$ is real, symmetric, and positive semi-definite [8]. This allows for the eigen-decomposition $\boldsymbol{L}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{\top}$, where $\boldsymbol{U}=[\mathbf{u}_{0},\ldots,\mathbf{u}_{n-1}]$ is an orthonormal matrix composed of eigenvectors $\mathbf{u}_{i}$, and eigenvalues $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{0},\ldots,\lambda_{n-1})$ are often interpreted as the graph frequencies.
Specifically, the eigenvectors above can be viewed as a set of spectral basis, allowing the Graph Fourier Transform (GFT) of the graph signal $\boldsymbol{z}$ to be expressed as:
@@
where $\hat{\boldsymbol{z}}\in\mathbb{R}^{n\times c}$ represents the spectral coefficients for different graph frequencies, with the square of each column indicating the “energy” of the corresponding frequency. Similarly, the inverse Graph Fourier Transform (iGFT) is then defined as:
@@
In our method, we associate graph vertices with key points in the point cloud, constructing the adjacency matrix $\boldsymbol{\mathcal{W}}$ considering the distances among these points.
Figure 2: (a) The overall of our PointGST. During the fine-tuning phase, we freeze the given pre-trained backbones and only fine-tune the proposed lightweight Point Cloud Spectral Adapter (PCSA). (b) The details of our proposed PCSA. We treat the point tokens as graph signals and then transfer the point tokens from the spatial domain to the spectral domain for tuning. The PCSA is injected into each transformer layer in a parallel paradigm.
@@ 4 Method
Fully fine-tuning a pre-trained point cloud model provides advanced performance, but substantial training costs accompany it. To address this challenge, we propose a novel method called PointGST, where the overall framework is illustrated in Fig. 2. Specifically, our method mainly consists of a regular pre-trained backbone and a series of lightweight Point Cloud Spectral Adapters (PCSA) integrated into each Transformer layer. During the fine-tuning, we keep the entire backbone frozen and only utilize a small number of trainable parameters (i.e., PCSA) to learn task-specific knowledge from the features of each instance. The PCSA transforms the point tokens from the spatial to the spectral domain, which de-correlates the inherent confusion of the point tokens from the frozen backbone. It efficiently fine-tunes fewer parameters for downstream tasks while introducing intrinsic information from the downstream point clouds for targeted tuning. As a result, it can significantly reduce the training cost and effectively connect the general knowledge from the pre-trained tasks with downstream tasks, leading to promising performance.
@@ 4.2 Efficient Fine-Tuning in Spectral Domain
  • Then, the proposed PCSA uses a shared linear layer to adapt the features. Following this, we recover the tokens to the spatial domain via inverse Graph Fourier Transform (iGFT) to align with the output of the transformer layer.
@@
Figure 3: (a) An intuitive visualization for our constructed graph, the adjacency matrix is symmetric, and the diagonal elements equal to 1. Smaller element values indicate greater geometric distances between point clouds. (b) The detailed process of our global and local spectral basis generation.
@@ 4.2.1 From Point Cloud to Graph
where $\min(\boldsymbol{\Delta})=\min(\delta_{i,j}),i\neq j$ represents the minimum value in $\boldsymbol{\Delta}$ excluding zero, which varies across different samples and is data-dependent. $\boldsymbol{I}\in\mathbb{R}^{n\times n}$ is an identity matrix. Using the geometric locations of key points and a data-dependent scaling strategy, we construct a graph with weights based on a distance metric between the points, incorporating all the desired properties. An intuitive visualization of our constructed graph for a point cloud is present in Fig. 3(a).
@@ 4.2.2 Multi-Scale Point Cloud Graphs
To comprehensively capture the underlying structure in point clouds, we generate a series of graphs that enable the calculation of both global and local eigenvectors, as shown in Fig. 3(b). Such an approach ensures a thorough usage of spectral features across different scales.
Global graph. We first treat all key points as an entirety and transform them into a global graph [^2] For simplicity, we directly utilize the adjacency matrix to represent a graph. $\boldsymbol{\mathcal{W}}_{\mathcal{G}}$. The details for translating the original point cloud to a graph are detailed in Sec. 4.2.1.
@@ 4.2.3 Point Cloud Spectral Adapter
Down-projection. Let us denote the input tokens as $\boldsymbol{T}_{in}\in\mathbb{R}^{n\times C}$, where $C$ is the input dimension. Here, we omit the class token and refer to $\boldsymbol{T}_{in}$ as the $n$ point tokens in this section. Our PCSA begins with a downward projection using trainable parameters $\boldsymbol{W}_{d}\in\mathbb{R}^{r\times C}$, with $r\ll C$. The low-rank graph signal $\boldsymbol{T}_{s}$ are obtained by:
@@
Spectral basis generation. After the down-projection, the graph signal $\boldsymbol{T}_{s}$ still remains in the spatial domain. Two preparations are required before applying Graph Fourier Transform (GFT) in the spatial point clouds, first to compute the Laplacian matrix $\boldsymbol{L}_{\mathcal{G}}\in\mathbb{R}^{n\times n}$ and $\boldsymbol{L}_{\mathcal{L}_{i}}\in\mathbb{R}^{m\times m}$ for the constructed the global graph $\boldsymbol{\mathcal{W}}_{\mathcal{G}}\in\mathbb{R}^{n\times n}$ and local sub-graphs $\boldsymbol{\mathcal{W}}_{\mathcal{L}_{i}}\in\mathbb{R}^{m\times m}$, respectively. We then perform eigenvalue decomposition to generate the spectral basis for the GFT and its inverse, as shown in Fig. 3(b).
@@
where the $\mathtt{act}$ refers to the Swish activation function [53]. We initialize the linear layer to zero during the initial stages for stable training.
@@
where both $\hat{\boldsymbol{T}}_{\mathcal{G}}^{f}$ and $\hat{\boldsymbol{T}}^{f}_{\mathcal{L}_{i}}$ are the converted spatial tokens. It should be noted that due to the introduced sorting and grouping strategy, the tokens of local spectral represent point cloud patches that do not align with the backbone part. Therefore, it is necessary to reorder the spatial local tokens to match the order of $\boldsymbol{T}_{s}$. Specifically, we concatenate the $\boldsymbol{T}_{\mathcal{L}_{i}}^{f}$ of all sub-graphs and reordered it, resulting in the tokens $\hat{\boldsymbol{T}}_{\mathcal{L}}^{f}\in\mathbb{R}^{n\times r}$:
@@
Up-projection and output. The PCSA finally employs a zero-initialized upward projection matrix $\boldsymbol{W}_{u}\in\mathbb{R}^{C\times r}$ and a manual scale $s\in\mathbb{R}$ to restore the fine-tuned compressed representations to the dimensions processed by the backbone:
@@
4.3 Theoretical Analysis
4.3.1 Insight of applying Graph Fourier Transform in PEFT
We summarize applying Graph Fourier Transform in PEFT into twofold: 1) The Graph Fourier Transform (GFT) conducts a signal decorrelation, resulting in a compact representation in the spectral domain. 2) Tuning in the spectral domain actually adjusts the “energy” of each graph frequency for the graph signal.
For discrete signals like point clouds, the graph Laplacian matrices can be considered as precision matrices within Gaussian-Markov Random Fields (GMRFs) [57]. This probabilistic perspective shows that the GFT approximates the Karhunen-Loève Transform (KLT) for signal decorrelation in GMRFs. Previous studies [12, 89, 90] also indicate that the GFT is an approximately optimal linear transform for de-correlating signals, which aids in geometric data compression. Since the GFT approximates the KLT for optimal signal decorrelation across various statistical processes, tuning in this domain is efficient and explicit.
Besides, our method for generating customized point graphs considers the relationship between points and adapts to the topology of point clouds, thereby injecting intrinsic knowledge of downstream tasks into the GFT process. By adopting the GFT, we can define the $\alpha_{i}=\mathbf{u}_{i}^{\top}\boldsymbol{T}$, where $\mathbf{u}_{i}\in\mathbb{R}^{1\times n}$ is the $i$-th eigenvector and $\boldsymbol{T}\in\mathbb{R}^{n\times r}$ is the graph signal. The term $|\alpha_{i}|^{2}$ is commonly referred to as the “energy” at the $i$-th graph frequency (corresponding to the $k$-th eigenvalue $\lambda_{i}$ of the Laplacian matrix) for signal $\boldsymbol{T}$. In other words, the shared linear layer in the Point Cloud Spectral Adapter (PCSA) adjusts $\alpha_{i}$ at each frequency, allowing our PointGST to implement a direct and effective fine-tuning paradigm.
Based on the above perceptions, we argue that the proposed PCSA can supplement the inner informed features from the fixed branch with explicit knowledge directly from point clouds for better feature integration. Specifically, the downward project $\boldsymbol{W}_{d}$ generates a low-ranked feature for each point patch, which acts as graph signals with key points as vertices. Additionally, the residual connection and zero-initialized strategy in our PCSA ensure training stability. The experiments in the following sections will validate the effectiveness and robustness of our architecture.
4.3.2 Training Cost
Our method requires only fine-tuning the proposed lightweight Point Cloud Spectral Adapter (PCSA) while freezing the entire pre-trained backbone. Specifically, the GFT and iGFT operations do not introduce extra parameters, while each Spectral Adapter module adds only $2\times r\times d$ from $\boldsymbol{W}_{d}$/$\boldsymbol{W}_{u}$, and $r\times r$ from the shared tuning Linear as trainable parameters (bias and layer normalization are ignored). The intermediate dimension $r$ is relatively small compared to $d$. For example, we only need to fine-tune about 0.6% of the parameters in the SOTA baseline, PointGPT-L [4], achieving promising results.
As the amount of downstream tasks or datasets requiring fine-tuning increases, we only need to store the task/dataset-specific PCSA and one original pre-trained models. As shown in Tab. I, when fine-tuning various downstream tasks, our PointGST significantly compresses model size compared to the baselines (Point-MAE [47] and PointGPT-L [4]).
We further analyze the training GPU memory, which is an important metric for evaluating the training cost of a neural network. As shown in Fig. 4, our method significantly reduces training memory as the batch size increases compared with Point-MAE [47]. In addition, compared with the previous representative PEFT methods for point cloud [88, 99], we present clear advantages, especially under the widely-used batch sizes such as 16 and 32.
TABLE I: The comparison of our PointGST and baseline methods in terms of storage usage when fine-tuning various downstream point cloud datasets or tasks.
Point-MAE [47] 22.1M 22.1M 22.1M 22.1M 22.1M 27.1M 27.0M 164.6M
PointGST (ours) 0.6M 0.6M 0.6M 0.6M 0.6M 5.6M 5.6M 14.2M
PointGPT-L [4] 360.5M 360.5M 360.5M 360.5M 360.5M 389.9M 389.9M 2582.3M
PointGST (ours) 2.4M 2.4M 2.4M 2.4M 2.4M 31.9M 31.9M 75.8M
Figure 4: The comparison of used training GPU memory.
5 Experiments Setup
5.1 Datasets and Implement Details
All the experiments are conducted on a single NVIDIA 3090 GPU. During fine-tuning, the parameters of the pre-trained backbones remain frozen, while only the newly added parameters are updated. The hyper-parameters $k$ and $r$ are set to 4 and 36, respectively. We employ a transpose format of Z space-filling curve [46] to scan the key points to construct the local graphs. This subsection introduces the implementation details along with the dataset.
ScanObjectNN [68] is a challenging 3D real-world objects dataset, consisting of $\sim$15151515K indoor object point cloud instances, across 15 categories. There are three widely used variants, including OBJ_BG, OBJ_ONLY, and PB_T50_RS, each with increasing complexity. The training setup is aligned with the compared baseline. Specifically, we employ the AdamW optimizer [41] with a weight decay of 0.05 and a cosine learning rate scheduler [42], starting with an initial learning rate of 5e-4 and a 10-epoch warm-up. The model is trained for 300 epochs with a batch size of 32. For consistency, we utilize 2048 points in the point cloud, divided into 128 groups of 32 points each.
ModelNet40 [79] is a commonly used synthesis 3D point cloud dataset that contains 12,311 CAD models of 40 categories. The network training settings follow ScanObjectNN. The input point cloud consists of 1,024 points, divided into 64 groups with 32 points per group.
ShapeNetpart [83] is a popular point-level synthetic object part segmentation benchmark that includes 16,881 instances across 16 object categories and 50 part categories. We use the AdamW optimizer with a weight decay of 1e-4 and the cosine learning rate scheduler, with an initial learning rate of 1e-4 and a 10-epoch warm-up. The models are trained for 300 epochs with a batch size of 16, using 2048 input points divided into 128 groups of 32 points each.
S3DIS [2] comprises six large-scale indoor areas, covering 273 million points from 13 categories. Following previous work [65], we advocate using Area 5 for evaluation purposes for better and fair generalization performance benchmarking. The cosine learning rate scheduler with an initial learning rate of 2e-4 is applied. We train our model for 60 epochs with a batch size of 32. Other configurations are the same as ShapeNetpart.

### v2

@@ 2.3 Spectral Methods for Point Clouds
Spectral-based methods have achieved significant progress in point cloud understanding [62, 63, 64, 65], either by modeling local details for fine-grained feature extraction [26, 66] or applying graph spectral convolutions for compact representations [62, 63, 64, 65]. For example, Wang et al.[67] combine spectral feature learning and clustering to address limitations of local aggregation, while Zhang et al.[68] estimate hypergraph spectral components in various settings. Other works include spectral domain attacks [69], efficient wavelet-based spectral transforms [66], and spectral-domain GANs for 3D shape generation [70].
Distinct from prior spectral approaches, this paper focuses on efficient parameter-efficient fine-tuning (PEFT). We propose PointGST, which fine-tunes pre-trained point cloud models in the spectral domain, differing in both motivation and technique from previous spectral-based methods.
@@ 3 Preliminary
This section revisits the parameter-efficient fine-tuning paradigm and the concept of Graph Fourier Transform.
3.1 Parameter-Efficient Fine-Tuning Paradigm
Given a downstream training set $\boldsymbol{\Gamma}\left(\boldsymbol{x};\boldsymbol{y}\right)$ ($\boldsymbol{x}$ is the input and $\boldsymbol{y}$ is the label), parameter-efficient fine-tuning (PEFT) provides a practical solution by adapting pre-trained models $\mathcal{F}$ to downstream tasks in a more efficient manner. Existing PEFT methods focus on tuning only a small subset of parameters $\boldsymbol{\theta}^{\prime}$. The optimized parameter $\boldsymbol{\theta}^{\prime*}$ can be represented as:
$\boldsymbol{\theta}^{\prime*}=\mathop{\arg\min}\limits_{\boldsymbol{\theta}^{\prime}}\ell\left(\mathcal{F}\left(\boldsymbol{x};\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right),\boldsymbol{y}\right),\left|\boldsymbol{\theta}^{\prime}\right|\ll\left|\boldsymbol{\theta}\right|,$ (1)
where $\ell$ represents the loss function for downstream tasks. The number of parameters in $\boldsymbol{\theta}^{\prime*}$, denoted $\left|\boldsymbol{\theta}^{\prime*}\right|$, equals $\left|\boldsymbol{\theta}^{\prime}\right|$ and is significantly smaller than the pre-trained $\boldsymbol{\theta}$. During fine-tuning, the pre-trained $\boldsymbol{\theta}$ remains fixed while only $\boldsymbol{\theta}^{\prime}$ is updated. Techniques for incorporating $\boldsymbol{\theta}^{\prime}$ include adding lightweight additional parameters and selecting a small subset of $\boldsymbol{\theta}$.
@@ 3.2 Graph Fourier Transform
Let us define a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\boldsymbol{\mathcal{W}}\}$, where $\mathcal{V}$ is the set of vertices ($|\mathcal{V}|=n$), $\mathcal{E}$ represents the edges, and $\boldsymbol{\mathcal{W}}$ is the adjacency matrix. In our context, a graph signal $\boldsymbol{z}\in\mathbb{R}^{n\times c}$ assigns a vector ($\boldsymbol{z}_{i}\in\mathbb{R}^{1\times c}$) to each vertex, while the entry $w_{i,j}$ in $\boldsymbol{\mathcal{W}}$ indicates the weight of the edge between vertices $i$ and $j$, reflecting their similarity. Then the Laplacian matrix [72] is defined as $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{\mathcal{W}}$, where $\boldsymbol{D}$ is a diagonal matrix with each element $d_{i,i}=\sum_{j=0}^{n-1}w_{i,j}$ representing the degree of vertex $i$. For an undirected graph with non-negative weights, $\boldsymbol{L}\in\mathbb{R}^{n\times n}$ is real, symmetric, and positive semi-definite [73]. This allows for the eigen-decomposition $\boldsymbol{L}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{\top}$, where $\boldsymbol{U}=[\mathbf{u}_{0},\ldots,\mathbf{u}_{n-1}]$ is an orthonormal matrix of eigenvectors $\mathbf{u}_{i}$, and the eigenvalues $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{0},\ldots,\lambda_{n-1})$ are often interpreted as the graph frequencies.
Specifically, the eigenvectors above can be viewed as a set of orthogonal spectral basis, allowing the Graph Fourier Transform (GFT) of the graph signal $\boldsymbol{z}$ to be expressed as:
@@
where $\hat{\boldsymbol{z}}\in\mathbb{R}^{n\times c}$ represents the spectral coefficients for different graph frequencies. Similarly, the inverse Graph Fourier Transform (iGFT) is then defined as:
@@
Our method associates graph vertices with key points in the point cloud, constructing the adjacency matrix $\boldsymbol{\mathcal{W}}$ considering the distances among these points.
Figure 2: (a) The overall of our PointGST. During the fine-tuning phase, we freeze the given pre-trained backbones and only fine-tune the proposed lightweight Point Cloud Spectral Adapter (PCSA). (b) The details of PCSA. We treat the point tokens as graph signals and then transfer the point tokens from the spatial domain to the spectral domain for tuning.
@@ 4 Method
Fully fine-tuning pre-trained point cloud models yields strong performance but incurs substantial training costs. To address this challenge, we propose a novel method called PointGST, where the overall framework is illustrated in Fig. 2. Specifically, our method mainly consists of an off-the-shelf pre-trained backbone and a series of lightweight Point Cloud Spectral Adapters (PCSA) integrated into each Transformer layer. During the fine-tuning, the entire backbone is frozen, and only a small set of trainable parameters (i.e., PCSA) are used to capture task-specific knowledge. The PCSA transforms the point tokens from the spatial to the spectral domain, which de-correlates the inherent confusion among the point tokens from the frozen backbone. It efficiently fine-tunes fewer parameters for downstream tasks while introducing intrinsic information from the downstream point clouds for targeted tuning. As a result, it can significantly reduce the training cost and effectively connect the general knowledge from the pre-training tasks with downstream tasks, leading to promising performance.
@@ 4.2 Efficient Fine-Tuning in Spectral Domain
  • Then, the PCSA uses a shared linear layer to adapt the features. Following this, we recover the tokens to the spatial domain via inverse Graph Fourier Transform (iGFT) to align with the output of the transformer layer.
@@ 4.2.1 From Point Cloud to Graph
where $\min(\boldsymbol{\Delta})=\min(\delta_{i,j},i\neq j)$ is the minimum value in $\boldsymbol{\Delta}$ excluding zero, which varies across different samples and is data-dependent. $\boldsymbol{I}\in\mathbb{R}^{n\times n}$ is an identity matrix. Using the geometric locations of key points and a data-dependent scaling strategy, we construct a graph with weights based on a distance metric between the points, incorporating all the desired properties.
Figure 3: The detailed process of global and local spectral basis generation.
@@ 4.2.2 Multi-Scale Point Cloud Graphs
To comprehensively capture the underlying structure in point clouds, we generate a series of graphs that enable the calculation of both global and local eigenvectors, as shown in Fig. 3. Such an approach ensures a thorough usage of spectral features across different scales.
Global graph. We first treat all key points as an entirety and transform them into a global graph [^2] For simplicity, we utilize the adjacency matrix to represent a graph. $\boldsymbol{\mathcal{W}}_{\mathcal{G}}$. The details for translating the original point cloud to a graph are detailed in Sec. 4.2.1.
@@ 4.2.3 Point Cloud Spectral Adapter
Down-projection. Let us denote the input tokens as $\boldsymbol{T}_{in}\in\mathbb{R}^{n\times C}$, where $C$ is the input dimension. Here, we omit the class token and refer to $\boldsymbol{T}_{in}$ as the $n$ point tokens in this section. Our PCSA begins with a downward projection using trainable parameters $\boldsymbol{W}_{d}\in\mathbb{R}^{r\times C}$, with $r\ll C$. The low-rank graph signal $\boldsymbol{T}_{s}$ is obtained by:
@@
Spectral basis generation. After the down-projection, the graph signal $\boldsymbol{T}_{s}$ still remains in the spatial domain. Two preparations are required before applying Graph Fourier Transform (GFT) in the spatial point clouds, first to compute the Laplacian matrix $\boldsymbol{L}_{\mathcal{G}}\in\mathbb{R}^{n\times n}$ and $\boldsymbol{L}_{\mathcal{L}_{i}}\in\mathbb{R}^{m\times m}$ for the constructed global graph $\boldsymbol{\mathcal{W}}_{\mathcal{G}}\in\mathbb{R}^{n\times n}$ and local sub-graphs $\boldsymbol{\mathcal{W}}_{\mathcal{L}_{i}}\in\mathbb{R}^{m\times m}$, respectively. We then perform eigenvalue decomposition to generate the spectral basis for the GFT and its inverse, as shown in Fig. 3.
@@
where the $\mathtt{act}$ refers to the Swish activation function [74]. We initialize the linear layer to zero for stable training.
@@
where both $\hat{\boldsymbol{T}}_{\mathcal{G}}^{f}$ and $\hat{\boldsymbol{T}}^{f}_{\mathcal{L}_{i}}$ are the converted spatial tokens. It should be noted that due to the introduced sorting and grouping strategy, the tokens of local spectral represent point cloud patches that do not align with the original token order in the backbone’s input. Therefore, it is necessary to reorder the spatial local tokens to match the order of $\boldsymbol{T}_{s}$. Specifically, we concatenate the $\boldsymbol{T}_{\mathcal{L}_{i}}^{f}$ of all sub-graphs and reordered it, resulting in the tokens $\hat{\boldsymbol{T}}_{\mathcal{L}}^{f}\in\mathbb{R}^{n\times r}$:
@@
Up-projection and output. The PCSA finally employs a zero-initialized up-projection matrix $\boldsymbol{W}_{u}\in\mathbb{R}^{C\times r}$ and a scale $s\in\mathbb{R}$ to restore the fine-tuned compressed representations to the dimensions processed by the backbone:
@@
Figure 4: An intuitive overview of spatial and spectral domains. (a) Tokens in the spatial domain after Point-MAE [8] pre-training, where red/blue indicate high/low cosine similarities to the zoomed-in area, respectively. (b) The spectral decomposition projects point tokens $\boldsymbol{T}$ on $n$ eigenvector of the Laplacian (using three as an example).
4.3 Insight of PEFT in the Spectral Domain
This section provides an intuitive understanding of the differences between spectral and spatial tuning and the insight of PEFT in the spectral domain.
The spatial domain tuning methods utilize tuning modules to process tokens representing localized patches of the point cloud, each corresponding to a specific region of points. As illustrated in Fig. 4(a), the zoomed-in area denotes one of these spatial tokens within a pre-trained model, where spatial fine-tuning adjusts the feature response at that patch. Besides, it can be found that even with similar geometric structures, features from the two plane wings of the point cloud show significant differences in output (reflected by their low cosine similarity). We argue that the pre-training tasks prioritize general point cloud representations without prior downstream tasks, resulting in the frozen model’s difficulty in distinguishing fine-grained point cloud structures in the spatial domain.
Our PointGST proposes the spectral domain tuning paradigm. We transfer the point tokens from the spatial domain into the spectral domain via the Graph Fourier Transform (GFT) theory. A toy example is shown in Fig. 4(b), we can decompose the original features $\boldsymbol{T}\in\mathbb{R}^{n\times r}$ into $n$ orthogonal Laplacian eigenvectors, where each eigenvector corresponds to a graph signal with a distinct level of smoothness. It is worth noting that the Laplace matrix is constructed based on the geometric attributes of each point cloud instance, and the resulting eigenvectors are shared across different Transformer layers, thus involving intrinsic information about the downstream point clouds. Then, the goal of spectral domain fine-tuning is to modify the projected coefficients or the position ($\cdots,\boldsymbol{T}_{i}^{f},\boldsymbol{T}_{j}^{f},\boldsymbol{T}_{k}^{f},\cdots$) of $\boldsymbol{T}$ within the $n$-dimensional eigenvector space.
In our context, the graph signal (compressed point tokens) $\boldsymbol{T}$ is defined on $n$ vertices. For simplicity, we assume $r=1$ here without loss of generality, as cases with $r>1$ can be reduced to multiple $r=1$ scenarios. Thus, Eq. 10 and Eq. 12 for $\boldsymbol{T}\in\mathbb{R}^{n\times 1}$ can be rewritten as:
$\boldsymbol{T}^{f}_{i}=\sum_{k=0}^{n-1}\boldsymbol{U}_{k,i}\boldsymbol{T}_{k}=\langle\mathbf{u}_{i},\boldsymbol{T}\rangle,~~\boldsymbol{T}=\sum_{i=0}^{n-1}\boldsymbol{T}^{f}_{i}\mathbf{u}_{i},$ (15)
where ${\boldsymbol{T}}^{f}_{i}\in\mathbb{R}$ indicates the projection of $\boldsymbol{T}$ onto the $i$-th graph Fourier basis vector $\mathbf{u}_{i}$ and also the $i$-th Fourier coefficient, as shown in Fig. 4(b). In the GFT, the Laplace matrix $\boldsymbol{L}$ functions similarly to the Laplace operator in the Fourier transform, using the eigenvectors of $\boldsymbol{L}$ as the basis for graph Fourier projection. $\boldsymbol{L}$ is defined as the difference between the degree matrix and the adjacency matrix, measuring variations of the graph through edge-based differentiation. We have the total variation defined as below:
$\boldsymbol{T}^{\top}\boldsymbol{L}\boldsymbol{T}={\boldsymbol{T}^{f}}^{\top}\boldsymbol{\Lambda}\boldsymbol{T}^{f}=\sum_{i=0}^{n-1}\lambda_{i}{\boldsymbol{T}^{f}_{i}}^{2}.$ (16)
From Eq. 16, it can be proved that to select a set of orthogonal graph signals with minimized total variances, those signals must be the eigenvectors of the Laplacian matrix [72]. Consequently, $\mathbf{u}_{0}$ corresponds to the smoothest graph signal and $\mathbf{u}_{n-1}$ to the least smooth, as they relate to the smallest eigenvalues $\lambda_{0}$ and the largest $\lambda_{n-1}$, respectively. Thus, the GFT decomposes a graph signal into components of varying smoothness, analogous to how the Fourier transform separates a function into different frequencies.
To summarize, intuitively, the GFT decomposes a graph signal (in our case, point tokens) into orthogonal graph signals of varying smoothness (eigenvectors of the Laplace matrix), enabling the original point token to be expressed as a linear combination of these signals. Using eigenvectors as a basis helps alleviate spatial domain confusion in the spectral domain.